OpenStudy (anonymous):

Solve the equation by applying the squaring property of equality. Be sure to check all solutions in the original equation. ( If there is no solution, enter NO SOLUTION.)

OpenStudy (anonymous):

\[\sqrt{2x+3}-3=4\] @charliealbert

OpenStudy (anonymous):

\[(\sqrt{2x+3})^{2}=7\] \[2x+3=7\] \[x=2\]

OpenStudy (anonymous):

2(2)+3-3=4 4+3-3=4 7-3=4 4=4!! Yes! THANK YOU! lol

OpenStudy (anonymous):

@Jack1 @whpalmer4 , this one ssaid incorrect :(

OpenStudy (jack1):

sqrt (2x+3)−3=4 sqrt (2x+3)−3+3 = 4+3 [sqrt (2x+3)]^2 = [7]^2 2x+3 = 49 2x+3-3 = 49-3 2x = 46 so x = 23... in my working anywhay...?

OpenStudy (anonymous):

O sorry, I forgot to elevate the seven XD

OpenStudy (anonymous):

It's ok lol